The market price for a bond that matures in 10 years is $ 900 and its even value is $ 1,000 and pays 8% interest (4% semi-annually). What is the expected rate of return on the bond? Without calculator.
Answer #1

To find the expected rate of return, we have to find the Yield to Maturity (YTM) of the bond. Although to calculate the exact YTM a calculator is required but for long dated bond the following formula gives estimate of Yield to maturity

YTM = (Annual Interest + ( Face value - Current Market Price)/(No of Years to Maturity))/((Face Value + Current Market Price)/2)

Data given in Question

Current Market Price: $900

Face Value: $1000

Annual Interest = $80 (Exact annual interest will be slightly higher as coupons are paid semiannually)

No of years to maturity: 10

So, using above formula and data given

YTM = (80 + ((1000 -900)/10))/(900+1000)/2 = (80 + 10)/(950) = 9.47%

So, expected rate of return is 9.47% approximately.